7) A 4.0 m length of gold wire is connected to a 1.5 V battery, and a current of 4.0 mA flows through it. What is the diameter of the wire? (The resistivity of gold is 2.44 × 10-8 Ω·m.) A) 17 µm B) 8.5 µm C) 48 µm D) 9.0 µm
Answer #1

R = rho*L/A

R = V/i

Equating both equations ,

1.5/4x10-3 = 2.44x10-8x4/pi*radius2

solving for radius , we get radius= 17 um